Factorise the expressions and divide them as directed.
12xy(9x^2 –16y^2 )÷4xy(3x+4y)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factorize 9x^2–16y^2, we have
9x^2–16y^2= (3x)^2–(4y)^2
= (3x+4y)(3x-4y) using identity: p2–q2= (p–q)(p+q)
Now, 12xy(9x^2–16y^2) ÷ 4xy(3x+4y)
= 12xy(3x+4y)(3x-4y) /4xy(3x+4y)
= 3(3x-4y)

Thanks